# Adaptive Parametric Activation

### v2

@@ Abstract
The activation function plays a crucial role in model optimisation, yet the optimal choice remains unclear. For example, the Sigmoid activation is the de-facto activation in balanced classification tasks, however, in imbalanced classification, it proves inappropriate due to bias towards frequent classes. In this work, we delve deeper in this phenomenon by performing a comprehensive statistical analysis in the classification and intermediate layers of both balanced and imbalanced networks and we empirically show that aligning the activation function with the data distribution, enhances the performance in both balanced and imbalanced tasks. To this end, we propose the Adaptive Parametric Activation (APA) function, a novel and versatile activation function that unifies most common activation functions under a single formula. APA can be applied in both intermediate layers and attention layers, significantly outperforming the state-of-the-art on several imbalanced benchmarks such as ImageNet-LT, iNaturalist2018, Places-LT, CIFAR100-LT and LVIS and balanced benchmarks such as ImageNet1K, COCO and V3DET. The code is available at https://github.com/kostas1515/AGLU.
Keywords:
Activation function Long-tailed learning
1 Introduction
@@
Image recognition has witnessed tremendous progress over the last years due to the use of deep learning, large image datasets such as ImageNet1K [17] and advances in model architectures [19, 28], learning algorithms [24, 69, 25], activation layers [31, 21, 27] and normalisation techniques [4, 40]. In this work, we focus on the activation layer of the network.
@@
In balanced image classification works, it was empirically shown that if the activation function is close to the real data distribution then the model converges faster because the learning objective becomes easier [27, 31]. Based on this, the GELU [31] and the PRELU [27] were proposed as alternatives to the commonly used RELU [21] and they were utilised inside the model’s layers to activate the intermediate features of the network. Similarly, in imbalanced image classification, many works have empirically shown that the Sigmoid or the Softmax activation functions, are inappropriate and using another activation function increases the performance [71, 34, 63, 1]. Based on that, the Gumbel activation [1] and the Balanced Softmax [71] activation were proposed and they were used inside the classification layer to predict the classes. In contrast to balanced classification, these works focused only on the classification layer and they disregarded the importance of the intermediate activations. To this end, there is no principled way of choosing the right activation function, and usually, practitioners use the best activation function, according to the task, through cross-validation or parameter-tuning.
@@
To this end, we empirically show that the commonly used Sigmoid activation function cannot generalise for both balanced and imbalanced learning, because it is non-parametric and does not align with the imbalanced data distribution. Motivated by this, we develop a novel Adaptive Parametric Activation (APA) function. APA allows the model to align its activations to both the balanced and imbalanced data distributions and reach great performance in both tasks. APA has several benefits, it unifies most previous activations functions such as the Sigmoid, the Gumbel [1], the RELU [21], the SiLU [31] and the GELU[31] under a common formula. Also, it uses two learnable parameters that allow the network to select the best activation function during optimisation enlarging the model’s capacity. Moreover, our APA is versatile, it can be used as a direct replacement to RELU, or it can replace the Sigmoid activation function inside the attention mechanism boosting the performance significantly and consistently. Finally, APA can be generalised to both imbalanced and balanced classification and detection tasks and surpass the state-of-the-art (SOTA). Our contributions are:
@@
  • We have validated the efficacy of APA on a range of long-tailed benchmarks, including ImageNet-LT [58], iNaturalist18 [84], Places-LT [58], CIFAR100-LT [8], LVIS [22] and balanced benchmarks such as ImageNet1k [17], COCO [55] and V3Det [88] largely surpassing the state of the art.
2 Preliminaries
Activation function. First, we show the importance of the activation function in model optimisation, following [1]. Let’s consider the example of binary classification, where $z$ is the input, $y\in\{0,1\}$ is the target class and $f(z)=W^{T}z+b$ is the classification network. The target class and the input $z$ are related as:
@@ 2 Preliminaries
$P(\bar{y}=1)=\frac{1}{1+\exp(-f(z))}=F_{\text{logistic}}(f(z);0,1)=1-F_{\text{logistic}}(-(f(z);0,1)$ (3)
Comparing Eq. 2 and Eq. 3, it is shown that when we use the Sigmoid, we assume that the error term $\epsilon$ follows the standard Logistic distribution. If we use Gumbel activation then we assume that the error follows the Gumbel distribution [1] and, in general, any activation function assumes a different distribution of $\epsilon$.
For this reason, the activation function can be seen as an initialisation point, or a prior belief of how the real data are distributed. If the prior is good, then the learning objective becomes easier and the performance is increased as it was shown empirically by many past studies [27, 1, 21, 77].
2.1 Balanced versus Imbalanced learning
@@
Classification logits. We train a MaskRCNN [26] ResNet50[28] on LVIS dataset [22], which is a highly imbalanced object detection dataset containing 1203 classes. After the model has converged, we perform inference and store the predicted classification logits, i.e., $f(z_{y})$ for every class $y$. Next, we perform histogram binning on the logits and we visualise the empirical distributions for the rare class puffin and the frequent class glove in Fig. 1 (b) and (c) respectively.
The rare class has a negative average logit value, because it is dominated by the frequent classes, which have higher average logit values as shown in (c). Regarding its distribution, it resembles the Gumbel distribution, because it has a heavy right tail and it is skewed. To quantify that, we calculate the statistical distance, using the Kolmogorov-Smirnov (KS) test [62], between all empirical class distributions and the Gumbel and Logistic theoretical distributions, shown in (a) and (e). As shown in (d), most classes have smaller distance to the Gumbel distribution. We repeat this test, using ResNet50 trained on the balanced ImageNet1K. As shown in (f) and (g), the classification logits are different this time, for example most average logit values are centered around zero and they are less skewed. In general, the logit distributions are closer to the logistic distribution as shown in (h). This explains why the Sigmoid activation achieves better performance in the balanced classification task and worse performance in imbalanced classification [1, 71, 34, 63]. Next, we show that data imbalance also affects the intermediate layers, by studying the channel attention as an example.
@@
Attention layer. Attention mechanism for input $X\in\mathbb{R}^{H\times W\times C}$ re-weights the input features $X$ by applying an attention function $A(X)$, i.e., $X^{\prime}=A(X)\otimes X$. For example, in Channel Attention (CA) [37], $X^{\prime}=\sigma(\text{MLP}(\text{GAP}_{c}(X)))\otimes X$, where $\text{GAP}_{c}\in\mathbb{R}^{1\times 1\times C}$ is Global Average Pooling and $\otimes$ is the element-wise product. In this case, the attention function is $A_{CA}(X)=\sigma(\text{MLP}(\text{GAP}_{c}(X)))$.
Figure 2: Visualisations of channel attention (A). In (a) the attention signals when training with imbalanced ImageNet and balanced ImageNet have similar variance in the first layer but completely different in the last most semantic layer in (b). The variance with ImageNet-LT training drops to zero for deeper layers as shown in (c), because the attention promotes only few frequent classes. In (d) and (e) the entropy of channel attention is smaller for the rare classes than the frequent classes in both i-Naturalist18 and ImageNet-LT training. In (f) and (g) the channel responses are smaller for the rare classes than the frequent ones for i-Naturalist18 and ImageNet-LT respectively.
Balanced vs Imbalanced channel attention. We train SE-ResNet50 [37] models (SE-R5) on balanced ImageNet-1K and imbalanced ImageNet-LT. After training the models, we analyse the average channel attention signals for a random batch of 128 test images. Fig. 2-a shows the output $A_{CA}$, in the first layer of SE-R50, Fig. 2-b shows the output, in the last layer, and Fig. 2-c shows the variance of $A_{CA}$ across all layers. As shown in Fig. 2-a, the attention is similar in the first layer and it is different in the last layer, as shown in Fig. 2-b. As displayed in Fig. 2-c, the attention variance with ImageNet1K (blue-curve) is larger and the attention reweights all channels and affects all classes. In contrast, for the imbalanced case (orange-curve), the attention signal has small variance, indicating that it is biased to some classes.
@@
$\begin{split}E_{l}=-\frac{1}{C}\sum_{i=1}^{C}[(A_{CA}(X_{l})\log(A_{CA}(X_{l}))+(1-A_{CA}(X_{l})\log(1-A_{CA}(X_{l}))]\end{split}$ (4)
@@
3 Method
As shown in the previous section, the degree of data imbalance affects both the classification logits and the intermediate layers. While it is possible to perform a statistical analysis and select the appropriate activation for the classification layer, this is difficult to do for all layers of the network, because first, the data distributions inside the layers dynamically change during training  [4, 40], and secondly, there is no one-to-one correspondence between classes and intermediate channels, which hinders attribution. For this reason, we propose the Adaptive Parametric Activation (APA):
@@ 3 Method
APA can adjust its activation rate, dynamically, according to the input’s distribution using two parameters $\kappa$ and $\lambda$, that can be learned during optimisation. $\kappa\in\mathbb{R}$ is the gain parameter that controls the function’s sensitivity. $\lambda\in(0,\infty)$ is the asymmetrical parameter that controls the function’s response rate to positive and negative inputs, allowing the model to have different learning degrees when the input is positive or negative. This function is also known as Richard’s curve [73] and it unifies the most common activation functions.
For example, if $\kappa=\lambda=1$ then APA becomes the Sigmoid activation that has a symmetric response rate for both positive and negative inputs and it is successful for balanced classification tasks. If $\kappa=1,\lambda\xrightarrow[]{}0$ then APA becomes the Gumbel activation that has an asymmetric response rate and it is successful for long-tailed instance segmentation tasks [1]. This behaviour is shown in Fig. 3 left and middle. Based on Eq. 5, we also define the Adaptive Generalised Linear Unit (AGLU):
$\begin{split}AGLU(z,\kappa,\lambda)=z\cdot\eta_{ad}(z,\kappa,\lambda)\\ \end{split}$ (6)
@@
AGLU has many interesting properties, for example, if $\kappa=\lambda=1$ then AGLU becomes the Sigmoid Linear Unit (SiLU) [31]. If $\kappa=1.702,\lambda=1$, then it becomes the Gaussian Error Linear Unit (GELU) [31]. If $\kappa\xrightarrow[]{}\infty$, then AGLU becomes ReLU[21] and if $\lambda\xrightarrow[]{}\infty$, then AGLU becomes the identity function, as shown in Fig. 3-right.
In other words, the $\kappa$ parameter controls the RELU-ness and the $\lambda$ parameter controls the leakage. Also, AGLU could be seen as a smoother version of PRELU, because $AGLU(z,1,\lambda\xrightarrow[]{}\infty)=PRELU(z,1)$. We compare $AGLU$ to most common activation functions, in more detail, in Table 1. The derivative of AGLU with respect to $\kappa$ is:
@@
Table 1: Comparison of different activation functions.
Name Formula Range
RELU [21] $\eta(z)=\max(0,z)$ $(0,\infty)$
Gaussian Unit [31] $\eta(z)=z\sigma(1.702z)$ $(-0.17,\infty)$
Sigmoid Unit [31] $\eta(z)=z\sigma(z)$ $(-0.28,\infty)$
Mish [64] $\eta(z)=z\tanh(\ln(1+\exp(z))$ $(-0.31,\infty)$
PRELU [27] $\eta(z,\kappa)=\max(0,z)+\kappa\min(0,z)$ (,)($\infty$,$\infty$)( - ∞ , ∞ )
ELU [11] $\eta(z,\kappa)=\max(0,z)+\kappa(\exp(\min(0,z))-1)$ $(-\kappa,\infty)$
AGLU (ours) $\eta(z,\kappa,\lambda)=z\cdot(\lambda\exp(-\kappa z)+1)^{\frac{1}{-\lambda}}$ (,)($\infty$,$\infty$)( - ∞ , ∞ )
In the end, our APA is versatile because it can be used not only as an error unit, replacing RELU, but also as an activation function inside the attention mechanism, replacing the Sigmoid activation.
4 Experiment Setup
Datasets. We use CIFAR100-LT [8] with exponential imbalance factor of 100 and 10, ImageNet-LT [58], Places-LT [58] and iNaturalist2018 [84] following the common long-tailed classification protocol. We report our results using top-1 accuracy on the balanced test sets, to fairly evaluate all classes. For ImageNet-LT, we split the categories according to their class frequency in the training set, into Many ($>$100 images), Medium (20$\sim$100 images) and Low ($<$20 images) and do per-group evaluation following [58]. Also, we use the LVISv1 [22] instance segmentation dataset, which has 100K training images and 1203 classes, that are grouped according to Frequent ($>$100 images), Common (10$\sim$100 images) and Rare ($<$10 images) classes. For this dataset, we report mask average precision $AP$, bounding box average precision $AP^{b}$ and $AP^{r}$, $AP^{c}$ and $AP^{f}$ which is mask average precision for rare, common and frequent classes respectively. Regarding balanced training, we perform experiments on ImageNet1K [17], COCO [55] and the recently proposed V3Det[88], which is a challenging large scale detection dataset with 13K classes and 243K images. We report top-1 accuracy for ImageNet1K, $AP^{b}$ and $AP^{m}$ for COCO and $AP^{b}$ for V3Det.
Table 2: Top-1 accuracy (%) on ImageNet-LT test set. E denotes ensemble.

### v3

@@ Abstract
The activation function plays a crucial role in model optimisation, yet the optimal choice remains unclear. For example, the Sigmoid activation is the de-facto activation in balanced classification tasks, however, in imbalanced classification, it proves inappropriate due to bias towards frequent classes. In this work, we delve deeper in this phenomenon by performing a comprehensive statistical analysis in the classification and intermediate layers of both balanced and imbalanced networks and we empirically show that aligning the activation function with the data distribution, enhances the performance in both balanced and imbalanced tasks. To this end, we propose the Adaptive Parametric Activation (APA) function, a novel and versatile activation function that unifies most common activation functions under a single formula. APA can be applied in both intermediate layers and attention layers, significantly outperforming the state-of-the-art on several imbalanced benchmarks such as ImageNet-LT, iNaturalist2018, Places-LT, CIFAR100-LT and LVIS. Also, we extend APA to a plethora of other tasks such as classification, detection, visual instruction following tasks, image generation and next-text-token prediction benchmarks. APA increases the performance in multiple benchmarks across various model architectures. The code is available at https://github.com/kostas1515/AGLU.
I Introduction
@@
Image recognition has witnessed tremendous progress over the last years due to the use of deep learning, large image datasets such as ImageNet1K [deng2009imagenet] and advances in model architectures [dosovitskiy2021an, he2016deep], learning algorithms [he2022masked, radford2021learning, he2020momentum], activation layers [hendrycks2016gaussian, glorot2011deep, he2015delving] and normalisation techniques [ba2016layer, ioffe2015batch]. In this work, we focus on the activation layer of the network.
@@
In balanced image classification works, it was empirically shown that if the activation function is close to the real data distribution then the model converges faster because the learning objective becomes easier [he2015delving, hendrycks2016gaussian]. Based on this, the GELU [hendrycks2016gaussian] and the PRELU [he2015delving] were proposed as alternatives to the commonly used RELU [glorot2011deep] and they were utilised inside the model’s layers to activate the intermediate features of the network. Similarly, in imbalanced image classification, many works have empirically shown that the Sigmoid or the Softmax activation functions, are inappropriate and using another activation function increases the performance [Ren2020balms, hong2021disentangling, menon2021longtail, alexandridis2022long]. Based on that, the Gumbel activation [alexandridis2022long] and the Balanced Softmax [Ren2020balms] activation were proposed and they were used inside the classification layer to predict the classes. In contrast to balanced classification, these works focused only on the classification layer and they disregarded the importance of the intermediate activations. To this end, there is no principled way of choosing the right activation function, and usually, practitioners use the best activation function, according to the task, through cross-validation or parameter-tuning.
@@
To this end, we empirically show that the commonly used Sigmoid activation function cannot generalise for both balanced and imbalanced learning, because it is non-parametric and does not align with the imbalanced data distribution. Motivated by this, we develop a novel Adaptive Parametric Activation (APA) function shown in Figure 1. APA allows the model to align its activations to both the balanced and imbalanced data distributions and reach great performance in both tasks. APA has several benefits, it unifies most previous activations functions such as the Sigmoid, the Gumbel [alexandridis2022long], the RELU [glorot2011deep], the SiLU [hendrycks2016gaussian] and the GELU[hendrycks2016gaussian] under a common formula. Also, it uses two learnable parameters that allow the network to select the best activation function during optimisation enlarging the model’s capacity. Moreover, our APA is versatile, it can be used as a direct replacement to RELU, or it can replace the Sigmoid activation function inside the attention mechanism boosting the performance significantly and consistently. Finally, APA can be generalised to both imbalanced and balanced tasks and surpass the state-of-the-art (SOTA). Our contributions are:
@@
  • We have validated the efficacy of APA on a range of long-tailed benchmarks, including ImageNet-LT [liu2019large], iNaturalist18 [van2018inaturalist], Places-LT [liu2019large], CIFAR100-LT [cao2019learning], LVIS [gupta2019lvis] largely surpassing the state of the art.
  • We comprehensively evaluate APA across various architectures and benchmarks, including image classification, object detection, VQA tasks, image generation and text classification and show that APA generalises to both large scale datasets and foundation models.
II Related Work
II-A Long-tailed image recognition.
Long-tailed image recognition can be grouped according to representation learning and classifier learning techniques. Representation learning techniques improve the feature quality through rare class features generation [vigneswaran2021feature, wang2021rsg, zang2021fasa], contrastive objectives [wang2021contrastive, cui2021parametric, zhu2022balanced, samuel2021distributional, li2022targeted, kang2021exploring], ensemble or fusion models [wang2021long, li2022nested, 9774921, zhou2020bbn, li2022trustworthy, cui2022reslt, cai2021ace], knowledge distillation [li2022nested, iscen2021cbd, he2021distilling, li2021self], knowledge transfer [liu2019large, parisot2022long, zhu2020inflated] and data augmentation [zhong2021improving, park2022majority, xu2021towards]. Classifier learning techniques enhance rare class classification through decoupled training [kang2019decoupling, zhang2021distribution, wang2020devil, kim2020adjusting, hsu2023abc], margin adjustment [menon2021longtail, Ren2020balms, hong2021disentangling, cao2019learning, hyun2022long, pan2021model, wang2021seesaw, wang2022c2am, zhao2022adaptive, alexandridis2023inverse, alexandridis2025fractal], cost-sensitive learning [cui2019class, khan2017cost, wang2017learning, huang2016learning], resampling [chawla2002smote, mahajan2018exploring, shen2016relay, zou2018unsupervised, gupta2019lvis, park2022majority, hong2022safa, zang2021fasa], dynamic loss adaptation based on batch statistics [hsieh2021droploss, tan2020equalization, wang2021adaptive], gradient statistics [tan2021equalization, li2022equalized], weight norms [wang2022c2am] and classification scores [feng2021exploring, he2022relieving]. These techniques are efficient, however they are applied either during the input phase or during the loss disregarding the intermediate activations.
II-B Attention Networks.
Attention networks, such as spatial attention [jaderberg2015spatial], channel attention [hu2018squeeze] and spatial-channel attention [woo2018cbam], are widely used in image recognition. Self-attention is the core mechanism of the Visual Transformer (ViT) architecture [dosovitskiy2021an], and recently many improvements have been made to the Transformer architecture [touvron2021training, touvron2021going, liu2021swin, xiao2021early] and its training procedure [steiner2021train, beyer2022better, touvron2022deit].
APA can be combined with many attention models like channel attention, Spatial, Spatial-Channel attention and self-attention models and further boost the performance.
II-C Activation functions
The RELU function [glorot2011deep] dominates the landscape of deep image classification, and it is especially used inside the convolutional networks, while the GELU and SiLU [hendrycks2016gaussian] are more commonly used inside the transformer network [dosovitskiy2021an] or the ConNext models [liu2022convnet, woo2023convnext]. The Gumbel activation is effective for imbalanced image recognition tasks [alexandridis2022long], and recently, it showed good performance for balanced tasks as well [das2025gompertz]. The PRELU function [he2015delving] is a generalisation of the RELU, because it linearly activates both positive and negative inputs and the ELU [clevert2015fast] is a follow-up work that enforces saturation on the negative inputs after some threshold. Compared to those, our APA, generalises the RELU, the GELU, SiLU and Gumbel and it can also be seen as a smoother version of PRELU. Recently, there has been increased interest in diffeomorphic activation functions used in image classification [chelly2024trainable], segmentation [freifeld2017transformations], parameter-efficient tuning [mantri2024digraf] and graph neural networks [mantri2024digraf]. However, these activation functions require extensive forward and backward computations and a user defined input-domain interval. In contrast, APA forward and backward is straightforward and it operates in the whole input domain.
III Preliminaries
III-A Activation function.
First, we show the importance of the activation function in model optimisation, following [alexandridis2022long]. Let’s consider the example of binary classification, where $z$ is the input, $y\in\{0,1\}$ is the target class and $f(z)=W^{T}z+b$ is the classification network. The target class and the input $z$ are related as:
@@ III Preliminaries
$P(\bar{y}=1)=\frac{1}{1+\exp(-f(z))}=F_{\text{logistic}}(f(z);0,1)$ (3)
$=1-F_{\text{logistic}}(-(f(z);0,1)$
Comparing Eq. 2 and Eq. 3, it is shown that when we use the Sigmoid, we assume that the error term $\epsilon$ follows the standard Logistic distribution. If we use Gumbel activation then we assume that the error follows the Gumbel distribution [alexandridis2022long] and, in general, any activation function assumes a different distribution of $\epsilon$.
For this reason, the activation function can be seen as an initialisation point, or a prior belief of how the real data are distributed. If the prior is good, then the learning objective becomes easier and the performance is increased as it was shown empirically by many past studies [he2015delving, alexandridis2022long, glorot2011deep, skorski2021revisiting].
III-B Balanced versus Imbalanced learning
@@
Classification logits. We train a MaskRCNN [he2017mask] ResNet50[he2016deep] on LVIS dataset [gupta2019lvis], which is a highly imbalanced object detection dataset containing 1203 classes. After the model has converged, we perform inference and store the predicted classification logits, i.e., $f(z_{y})$ for every class $y$. Next, we perform histogram binning on the logits and we visualise the empirical distributions for the rare class puffin and the frequent class glove in Fig. 2 (b) and (c) respectively.
The rare class has a negative average logit value, because it is dominated by the frequent classes, which have higher average logit values as shown in (c). Regarding its distribution, it resembles the Gumbel distribution, because it has a heavy right tail and it is skewed. To quantify that, we calculate the statistical distance, using the Kolmogorov-Smirnov (KS) test [massey1951kolmogorov], between all empirical class distributions and the Gumbel and Logistic theoretical distributions, shown in (a) and (e). As shown in (d), most classes have smaller distance to the Gumbel distribution. We repeat this test, using ResNet50 trained on the balanced ImageNet1K. As shown in (f) and (g), the classification logits are different this time, for example most average logit values are centered around zero and they are less skewed. In general, the logit distributions are closer to the logistic distribution as shown in (h). This explains why the Sigmoid activation achieves better performance in the balanced classification task and worse performance in imbalanced classification [alexandridis2022long, Ren2020balms, hong2021disentangling, menon2021longtail]. Next, we show that data imbalance also affects the intermediate layers, by studying the channel attention as an example.
@@
Attention layer. Attention mechanism for input $X\in\mathbb{R}^{H\times W\times C}$ re-weights the input features $X$ by applying an attention function $A(X)$, i.e., $X^{\prime}=A(X)\otimes X$. For example, in Channel Attention (CA) [hu2018squeeze], $X^{\prime}=\sigma(\text{MLP}(\text{GAP}_{c}(X)))\otimes X$, where $\text{GAP}_{c}\in\mathbb{R}^{1\times 1\times C}$ is Global Average Pooling and $\otimes$ is the element-wise product. In this case, the attention function is $A_{CA}(X)=\sigma(\text{MLP}(\text{GAP}_{c}(X)))$.
Figure 3: Visualisations of channel attention (A). In (a) the attention signals when training with imbalanced ImageNet and balanced ImageNet have similar variance in the first layer but completely different in the last most semantic layer in (b). The variance with ImageNet-LT training drops to zero for deeper layers as shown in (c), because the attention promotes only the frequent classes. In (d) and (e) the entropy of channel attention is smaller for the rare classes than the frequent classes in both i-Naturalist18 and ImageNet-LT training. In (f) and (g) the channel responses are smaller for the rare classes forboth i-Naturalist18 and ImageNet-LT. These observations shows that channel attention is not robust during imbalanced training, motivating the development of APA.
Balanced vs Imbalanced channel attention. We train SE-ResNet50 [hu2018squeeze] models (SE-R50) on balanced ImageNet-1K and imbalanced ImageNet-LT. After training the models, we analyse the average channel attention signals for a random batch of 128 test images. Fig. 3-a shows the output $A_{CA}$, in the first layer of SE-R50, Fig. 3-b shows the output, in the last layer, and Fig. 3-c shows the variance of $A_{CA}$ across all layers. As shown in Fig. 3-a, the attention is similar in the first layer and it is different in the last layer, as shown in Fig. 3-b. As displayed in Fig. 3-c, the attention variance with ImageNet1K (blue-curve) is larger and the attention reweights all channels and affects all classes. In contrast, for the imbalanced case (orange-curve), the attention signal has small variance, indicating that it is biased to some classes.
@@
$\begin{split}E_{l}=-\frac{1}{C}\sum_{i=1}^{C}[(A_{CA}(X_{l})\log(A_{CA}(X_{l}))\\ +(1-A_{CA}(X_{l})\log(1-A_{CA}(X_{l}))]\end{split}$ (4)
@@
IV Method
IV-A Adaptive Parametric Activation
As shown in the previous section, the degree of data imbalance affects both the classification logits and the intermediate layers. While it is possible to perform a statistical analysis and select the appropriate activation for the classification layer, this is difficult to do for all layers of the network, because first, the data distributions inside the layers dynamically change during training  [ba2016layer, ioffe2015batch], and secondly, there is no one-to-one correspondence between classes and intermediate channels, which hinders attribution. For this reason, we propose the Adaptive Parametric Activation (APA):
@@ IV-A Adaptive Parametric Activation
APA can adjust its activation rate, dynamically, according to the input’s distribution using two parameters $\kappa$ and $\lambda$, that can be learned during optimisation. $\kappa\in\mathbb{R}$ is the gain parameter that controls the function’s sensitivity. $\lambda\in(0,\infty)$ is the asymmetrical parameter that controls the function’s response rate to positive and negative inputs, allowing the model to have different learning degrees when the input is positive or negative. This function is also known as Richard’s curve [richards1959flexible] and it unifies the most common activation functions.
For example, if $\kappa=\lambda=1$ then APA becomes the Sigmoid activation that has a symmetric response rate for both positive and negative inputs and it is successful for balanced classification tasks. If $\kappa=1,\lambda\xrightarrow[]{}0$ then APA becomes the Gumbel activation that has an asymmetric response rate and it is successful for long-tailed instance segmentation tasks [alexandridis2022long]. This behaviour is shown in Fig. 1 left and middle. Based on Eq. 5, we also define the Adaptive Generalised Linear Unit (AGLU):
[…]
@@
AGLU has many interesting properties, for example, if $\kappa=\lambda=1$ then AGLU becomes the Sigmoid Linear Unit (SiLU) [hendrycks2016gaussian]. If $\kappa=1.702,\lambda=1$, then it becomes the Gaussian Error Linear Unit (GELU) [hendrycks2016gaussian]. If $\kappa\xrightarrow$$\infty$, then AGLU becomes ReLU[glorot2011deep] and if $\lambda\xrightarrow$$\infty$, then AGLU becomes the identity function, as shown in Fig. 1-right.
In other words, the $\kappa$ parameter controls the RELU-ness and the $\lambda$ parameter controls the leakage. Also, AGLU could be seen as a smoother version of PRELU, because $AGLU(z,1,\lambda\xrightarrow[]{}\infty)=PRELU(z,1)$. We compare AGLU to most common activation functions, in more detail, in Table I. The derivative of AGLU with respect to $\kappa$ is:
@@
TABLE I: Comparison of different activation functions.
Name Formula Output Range
RELU [glorot2011deep] $\eta(z)=\max(0,z)$ $(0,\infty)$
Gaussian Unit [hendrycks2016gaussian] $\eta(z)=z\sigma(1.702z)$ $(-0.17,\infty)$
Sigmoid Unit [hendrycks2016gaussian] $\eta(z)=z\sigma(z)$ $(-0.28,\infty)$
Mish [misra2019mish] $\eta(z)=z\tanh(\ln(1+\exp(z))$ $(-0.31,\infty)$
PRELU [he2015delving] $\eta(z,\kappa)=\max(0,z)+\kappa\min(0,z)$ $(-\infty,\infty)$
ELU [clevert2015fast] $\eta(z,\kappa)=\max(0,z)+\kappa(\exp(\min(0,z))-1)$ $(-\kappa,\infty)$
AGLU (ours) $\eta(z,\kappa,\lambda)=z\cdot(\lambda\exp(-\kappa z)+1)^{\frac{1}{-\lambda}}$ $(-\infty,\infty)$
APA is versatile and it has many use cases. For example it can be uses in the intermediate layers, i.e. replacing RELU, or inside the attention mechanism, replacing the Sigmoid activation.
IV-B Non-Linearity and Adaptability
APA ’s parameters control the amount of non-linearity added to the model. To measure this, we follow the newly proposed Neural Redshift framework [teney2024neural] and use the affinity score $\rho_{aff}$ [bouniot2025alexnet] to study the behavior of a randomly initialised three hidden-layer MLP with a scalar output. We initialise the MLP weights using various schemes such as Xavier [glorot2010understanding] and Kaiming [he2015delving] and the MLP biases with the uniform distribution U(1,1)U(-1,1), we feed an input of $64^{2}$ evenly spaced 2D coordinates on the [1,1][-1,1] grid and visualise the output in grayscale in Figure 5.
If RELU is used as nonlinearity, then the MLP has the biggest amount of non-linearity, because it produces the smallest affinity scores $\rho_{aff}$ and the most complex outputs for all the initialisation schemes as shown in the first column. If AGLU with small $\lambda$ and larger $\kappa$ values is used then the AGLU-MLP becomes similar to the RELU-MLP, i.e. it is more non-linear and it has smaller affinity scores $\rho_{aff}$ and more complex outputs as shown in the the second column. Decreasing the $\kappa$ or increasing the $\lambda$ values of AGLU makes the MLP less non-linear, i.e. the affinity scores increase and the outputs become smoother as shown in the third and fourth column respectively.
This has two effects, it allows the users to smoothly control the inductive bias of the model, i.e. by fixing the amount of non-linearity or it enables to learn the best nonlinearities by optimization.
Figure 5: Visualisation of the output of a random MLP, initialised with Xavier-Uniform, Xavier-Normal, Kaiming-Uniform and Kaiming-Normal, across the rows, and $RELU(z)$, $AGLU(z,2,0)$, $AGLU(z,1,0)$ and $AGLU(z,1,1)$ across the columns. The RELU MLP is the most non-linear, for all weight initialisations, because it has the smallest affinity scores $\rho_{aff}$ and most complex outputs as shown in the first column. Increasing the AGLU’s $\kappa$ value lowers $\rho_{aff}$ and makes the MLP more non-linear, as shown in the second column, while decreasing the $\kappa$ or increasing the $\lambda$ values increases $\rho_{aff}$ and makes it more linear as shown in the third and forth columns respectively.
V Experiments in Long-tail learning
V-A Datasets.
[…]
TABLE II: Top-1 accuracy (%) on ImageNet-LT test set. E denotes ensemble.